### v2

Abstract
We present PlainMamba: a simple non-hierarchical state space model (SSM) designed for general visual recognition. The recent Mamba model has shown how SSMs can be highly competitive with other architectures on sequential data and initial attempts have been made to apply it to images. In this paper, we further adapt the selective scanning process of Mamba to the visual domain, enhancing its ability to learn features from two-dimensional images by (i) a continuous 2D scanning process that improves spatial continuity by ensuring adjacency of tokens in the scanning sequence, and (ii) direction-aware updating which enables the model to discern the spatial relations of tokens by encoding directional information. Our architecture is designed to be easy to use and easy to scale, formed by stacking identical PlainMamba blocks, resulting in a model with constant width throughout all layers. The architecture is further simplified by removing the need for special tokens. We evaluate PlainMamba on a variety of visual recognition tasks, achieving performance gains over previous non-hierarchical models and is competitive with hierarchical alternatives. For tasks requiring high-resolution inputs, in particular, PlainMamba requires much less computing while maintaining high performance. Code and models are available at: https://github.com/ChenhongyiYang/PlainMamba.
1 Introduction
Developing high-performing visual encoders has always been one of the most important goals in computer vision [He et al.(2016)He, Zhang, Ren, and Sun, Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He, Simonyan and Zisserman(2015), Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby, Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo]. With high-quality visual features, a broad range of downstream tasks, such as semantic segmentation [Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba, Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Xiao et al.(2018)Xiao, Liu, Zhou, Jiang, and Sun, Wang et al.(2023)Wang, Kim, Feris, and Betke], object recognition [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, He et al.(2016)He, Zhang, Ren, and Sun, Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He, Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] and detection [Lin et al.(2017b)Lin, Goyal, Girshick, He, and Dollar, He et al.(2017)He, Gkioxari, Dollar, and Girshick, Ren et al.(2015)Ren, He, Girshick, and Sun] can be tackled with relative ease. Early methods for extracting visual representations relied on hand-crafted features such as SIFT  [Lowe(2004)] and SURF [Bay et al.(2006)Bay, Tuytelaars, and Van Gool]. A big breakthrough then came with the adoption of convolutional neural networks (CNNs) that process images with local contexts and enforce spatial equivariance [Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton, He et al.(2016)He, Zhang, Ren, and Sun, Simonyan and Zisserman(2015)]. Recently, vision transformers (ViTs) [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] obviated the need for such enforced inductive biases in favour of learnable contexts that operate on image patches [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin, Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo]. However, despite the overwhelming success of transformers and their self-attention mechanism [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova, Brown et al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei, Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], the quadratic cost of attention has proved to be an obstacle to further scaling such models.
This has invigorated interest in state space models (SSMs) [Williams et al.(2007)Williams, Lawrence, et al., Gu and Dao(2023), Gu et al.(2023)Gu, Johnson, Timalsina, Rudra, and Ré, Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. Due to their close ties to linear recurrent networks, SSMs have the benefits of potentially infinite context lengths while maintaining linear complexity in the input sequence length [Gu and Dao(2023)], offering substantial speedups compared to attention. However, it took several notable advances to make SSMs effective at learning competitive representations, including enforcing state space variables to be orthogonal basis projections [Gu et al.(2023)Gu, Johnson, Timalsina, Rudra, and Ré]. The recent Mamba [Gu and Dao(2023)] architecture further aligned SSM-based models with modern transformers, such as making the state space variables input-dependent — much like queries, keys and values in self-attention. When being used for NLP, those designs led to a state space model that could scale to the sizes and performances of modern transformer-based LLMs [Brown et al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei, Touvron et al.(2023)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al.], while improving inference effiency.
Figure 1: While hierarchical visual encoders may demonstrate superior accuracy on open-source visual recognition benchmarks, the plain non-hierarchical models have had more widespread use because of their simple structure. We investigate the potential of the plain Mamba model in visual recognition.
There is now understandable interest in adapting the Mamba architecture to the visual domain [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Li et al.(2024)Li, Singh, and Grover, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. However, before we start doing that, we need to think about under what guidelines should we design our new model. As we show in Figure 1, by examining the development of recently proposed visual encoders, we find that adding more inductive biases, e.g., hierarchical structure, to the plain model such as DeiT can indeed improve a model’s performance on open-source benchmarks like ImageNet. However, we should not ignore the fact that the plain ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] is widely used by several popular vision foundation models [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick, Liu et al.(2023b)Liu, Li, Wu, and Lee, Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski, et al(2022)], which suggests that simplicity in architecture design is key for multiple reasons. Firstly, maintaining a constant model width (i.e. non-hierarchical) makes it much easier to integrate features from multiple levels, as is common in dense prediction tasks such as semantic segmentation [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick]. It also becomes easier to combine features across different modalities such as in CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] or LLaVa [Liu et al.(2023b)Liu, Li, Wu, and Lee] or as parts of increasingly complex AI-powered systems. Furthermore, simpler components can be more easily optimized for hardware acceleration [Dao et al.(2022a)Dao, Fu, Ermon, Rudra, and Ré]. In addition, it has also been observed that the over-crafted models may lead to a significant gap between the performance on commonly used benchmarks and downstream tasks [Ericsson et al.(2021)Ericsson, Gouk, and Hospedales, Beyer et al.(2020)Beyer, Hénaff, Kolesnikov, Zhai, and Oord]. This means benchmark performance may no longer reflect real-world usefulness, as over-engineering tends to increase model complexity and thus make it harder for others to re-use.
Motivated by the above findings, we propose PlainMamba: a simple Mamba architecture for visual recognition. This model integrates ideas from CNNs, Transformers and novel SSM-based models with an aim to providing easy-to-use models for the vision modality. Compared to previous visual state space models [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], we simplify the architecture by maintaining constant model width across all layers of the network via stacking identical blocks as well as removing the need for CLS tokens. This allows for easy scaling and model re-use, while achieving competitive performances.
Our contributions are as follows: (1) We propose a new visual state space model we call PlainMamba. This architecture improves and simplifies previous attempts at extending the Mamba architecture to the visual modality. (3) We improve the SSM block by adapting selective scanning to better process 2D spatial inputs, in two ways. (i) Our continuous 2D scanning approach ensures that the scanning sequence is spatially continuous to improve semantic continuity. (ii) Our direction-aware updating, inspired by positional encoding, allows the model to encode the directionality of each scanning order to further improve spatial context. (3)We test our PlainMamba architecture using three different sizes (7M, 26M and 50M) and show how they perform competitively on a range of tasks, from ImageNet1K classification to semantic segmentation and object detection. Specifically, we show that PlainMamba outperforms its non-heretical counterparts, including SSMs and Transformers, while performing on par with the hierarchical competitors.
2 Related Work
Visual Feature Extractors. How to effectively extract visual features from images has been a long-standing challenge in computer vision. In the early years of deep learning, CNNs [Simonyan and Zisserman(2015), Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton, He et al.(2016)He, Zhang, Ren, and Sun, Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He] dominated the model architecture landscape. Their induced spatial prior, through the use of convolutional filters, exploits the locality of visual features. Furthermore, stacking multiple layers increases their receptive field. Many different CNN backbone architectures have been proposed over the years [Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton, Huang et al.(2017)Huang, Liu, van der Maaten, and Weinberger, Chollet(2016)], introducing new ways of exploiting spatial information [Simonyan and Zisserman(2015), Xie et al.(2016)Xie, Girshick, Dollár, Tu, and He], building deeper models [He et al.(2016)He, Zhang, Ren, and Sun, Szegedy et al.(2015)Szegedy, Liu, Jia, Sermanet, Reed, Anguelov, Erhan, Vanhoucke, and Rabinovich], improving efficiency [Sandler et al.(2018)Sandler, Howard, Zhu, Zhmoginov, and Chen, Tan and Le(2019), Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár], adding multi-scale connections [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox], scaling architectures [Woo et al.(2023)Woo, Debnath, Hu, Chen, Liu, Kweon, and Xie], and introducing attention mechanisms [Wang et al.(2017)Wang, Jiang, Qian, Yang, Li, Zhang, Wang, and Tang, Bello(2021), Srinivas et al.(2021)Srinivas, Lin, Parmar, Shlens, Abbeel, and Vaswani, Wang et al.(2020)Wang, Wu, Zhu, Li, Zuo, and Hu, Hu et al.(2018)Hu, Shen, and Sun, Chen et al.(2019)Chen, Fan, Xu, Yan, Kalantidis, Rohrbach, Yan, and Feng]. In recent years, ViTs have become a powerful tool for image modeling [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]. Compared to CNNs, they make fewer assumptions about data (feature locality [Xu et al.(2021b)Xu, Zhang, Zhang, and Tao], translation and scale invariance). By replacing the convolutional layers with self-attention modules, transformers can capture global relationships and have achieved state-of-the-art results on many common image benchmarks [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei, Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick, Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba]. To adapt the original transformer architecture [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] for vision tasks, images are split into patches and converted into tokens before being fed into the transformer encoder. Within this framework, numerous works have focused on pushing the performance (e.g. LeViT, [Graham et al.(2021)Graham, El-Nouby, Touvron, Stock, Joulin, Jégou, and Douze] combining transformer encoder layers and convolutions), or on reducing the costly quadratic complexity of self-attention [Dao et al.(2022b)Dao, Fu, Ermon, Rudra, and Ré, Dao(2024)]. Another popular extension to ViT architectures has been the addition of hierarchical structures [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Xu et al.(2021a)Xu, Xu, Chang, and Tu, Wu et al.(2021)Wu, Xiao, Codella, Liu, Dai, Yuan, and Zhang, Fan et al.(2021)Fan, Xiong, Mangalam, Li, Yan, Malik, and Feichtenhofer], similar to the multi-scale feature pyramids used in CNNs. The Swin Transformer [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], for instance, uses shifted windows to share feature information across scales. These multi-scale features are then used for a wide range of downstream tasks. Recent research has explored ways of using these hierarchical features within ViTs themselves [Ren et al.(2022)Ren, Zhou, He, Feng, and Wang, Guo et al.(2022)Guo, Han, Wu, Tang, Chen, Wang, and Xu, Li et al.(2022b)Li, Wu, Fan, Mangalam, Xiong, Malik, and Feichtenhofer, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo, Hatamizadeh et al.(2022)Hatamizadeh, Yin, Kautz, and Molchanov, Chen et al.(2022a)Chen, Panda, and Fan, d’Ascoli et al.(2021)d’Ascoli, Touvron, Leavitt, Morcos, Biroli, and Sagun, Lee et al.(2022)Lee, Kim, Willette, and Hwang, Shi(2024)]. Some works [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] have examined the use of multi-resolution features as attention keys and values to learn multi-scale information. However, these extensions add complexity to the model and make it harder to effectively use its features in later stages, thus hindering widespread adoption. Indeed, recent works [Li et al.(2022a)Li, Mao, Girshick, and He, Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang] return to the original ViT architecture, as its non-hierarchical nature greatly simplifies the use of its features. In particular, the plain ViT provides greater flexibility for pre-training and fine-tuning on different tasks.
Figure 2: (a) The overall architecture of the proposed PlainMamba. PlainMamba does not have a hierarchical structure, it instead stacks $L$ identical PlainMamba block to form the main network. For image classification, it uses global average pooling instead of the CLS to gather global information. (b) Architecture of PlainMamba block, which is similar to the Mamba [Gu and Dao(2023)] block where the selective scanning is combined with a gated MLP. (c) The proposed Direction-Aware Updating, where a series of learnable parameters $\mathbf{\Theta}_{k}$ are combined with the data-dependent updating parameters to explicitly inject relative 2D positional information into the selective scanning process.
State Space Models. State Space Models (SSMs) have emerged as efficient alternatives to transformers and CNNs due to their ability to scale linearly with sequence length [Gu et al.(2021)Gu, Johnson, Goel, Saab, Dao, Rudra, and Re, Sun et al.(2023)Sun, Dong, Huang, Ma, Xia, Xue, Wang, and Wei]. SSMs transform the state space to effectively capture dependencies over extended sequences. To alleviate the initial computational cost of such models, S4 [Gu et al.(2022)Gu, Goel, and Re] enforced low-rank constraints on the state matrix and S5 [Smith et al.(2023)Smith, Warrington, and Linderman] introduced parallel scanning to further improve efficiency. Furthermore, H3 [Fu et al.(2023)Fu, Dao, Saab, Thomas, Rudra, and Ré] achieved competitive results on common benchmarks by improving the hardware utilization. Lastly, Mamba [Gu and Dao(2023)] parameterized the SSM matrices as functions of the input, thus allowing it to act as a learnable selection mechanism and providing greater flexibility. Follow-up works have extended selective SSMs for images [Baron et al.(2023)Baron, Zimerman, and Wolf, Ruan and Xiang(2024), Liu et al.(2024a)Liu, Yang, Zhou, Xi, Yu, Yu, Liang, Shi, Zhang, Zheng, et al., Wang et al.(2024)Wang, Zheng, Zhang, Cui, and Li, Wang and Ma(2024), Ma et al.(2024)Ma, Li, and Wang, Pei et al.(2024)Pei, Huang, and Xu, Liu et al.(2024b)Liu, Yang, Zhou, Xi, Yu, Yu, Liang, Shi, Zhang, Zheng, et al.] and videos [Nguyen et al.(2022)Nguyen, Goel, Gu, Downs, Shah, Dao, Baccus, and Ré] using a hierarchical structure [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] and bidirectional blocks [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], while Mamba-ND [Li et al.(2024)Li, Singh, and Grover] introduces an architecture for multi-dimensional data. MambaIR [Guo et al.(2024)Guo, Li, Dai, Ouyang, Ren, and Xia] tackles image restoration, and Pan-Mamba [He et al.(2024)He, Cao, Yan, Li, Xie, Zhang, and Zhou] works on pan-sharpening. DiS [Fei et al.(2024)Fei, Fan, Yu, and Huang] introduces SSMs to diffusion models by replacing the U-Net with an SSM backbone. While drawing inspiration from the above works, PlainMamba improves Mamba’s [Gu and Dao(2023)] selective SSM block by adding wider depth-wise convolutions. In contrast to the Cross-Scan Module (CSM) [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] and Mamba-ND [Li et al.(2024)Li, Singh, and Grover], PlainMamba respects the spatio-sequential nature of image patches (see Figure 2). As opposed to [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], we do not use the CLS token.
Simplifying Visual Feature Extractors. Simplifying and unifying existing methods is equally important as improving performance. Plain architectures are robust, conceptually simpler, and scale better. ViTs [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] remove the pyramid structure of CNNs by converting images into patched tokens. This way, they easily adapt the transformer architecture for visual tasks. Another trick that stems form sequence modeling is the usage of CLS tokens for prediction, which have proven to be unnecessary for visual tasks [Zhai et al.(2022)Zhai, Kolesnikov, Houlsby, and Beyer]. FlexiVit [Beyer et al.(2023)Beyer, Izmailov, Kolesnikov, Caron, Kornblith, Zhai, Minderer, Tschannen, Alabdulmohsin, and Pavetic] unified into a single architecture images with different input resolutions, and GPViT [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang] improved feature resolution with a non-hierarchical transformer. Similarly, ConvNext [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] introduced a simple CNN model that competed with state-of-the-art transformer methods. Other works, like MLP-Mixer [Tolstikhin et al.(2021)Tolstikhin, Houlsby, Kolesnikov, Beyer, Zhai, Unterthiner, Yung, Steiner, Keysers, Uszkoreit, et al.] and follow-up works [Hou et al.(2022)Hou, Jiang, Yuan, Cheng, Yan, and Feng], have introduced simple architectures using only multi-layer perceptrons. The plain non-hierarchical ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] has served as a simple building block for many diverse tasks. SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] uses a pre-trained ViT as image encoder with minimal changes for image segmentation at large scale. DinoV2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski, Darcet et al.(2023)Darcet, Oquab, Mairal, and Bojanowski] uses a ViT to learn general-purpose visual features by pretraining models on curated datasets with self-supervision. Similarly, the image encoder for the CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] model consists of a basic ViT with minor modifications, allowing image-text representations to be learned with a contrastive objective. DALLE-2 [et al(2022)] incorporates a ViT image encoder to extract visual features that are used for text-conditional image generation. LlaVA [Liu et al.(2023b)Liu, Li, Wu, and Lee, Liu et al.(2023a)Liu, Li, Li, and Lee] combines a vision encoder (pretrained ViT from CLIP) and an LLM for vision-language tasks.
3 Method
3.1 Preliminaries
State Space Models. SSMs are typically used to model a continuous linear time-invariant (LTI) system [Williams et al.(2007)Williams, Lawrence, et al.] where an input signal $x(t)\in\mathbb{R}$ is mapped to its output signal $y(t)\in\mathbb{R}$ through a state variable $h(t)\in\mathbb{R}^{m}$ with the following rules:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)\mathrm{,}~{}~{}~{}y(t)=\mathbf{C}h^{\prime}(t)+\mathbf{D}x(t)$ (1)
where $\mathbf{A}\in\mathbb{R}^{m\times m}$, $\mathbf{B}\in\mathbb{R}^{m\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times m}$ and $\mathbf{D}\in\mathbb{R}^{1\times 1}$ are parameters. To make the above system usable for a discrete system, e.g., a sequence-to-sequence task, a timescale parameter $\mathbf{\Delta}$ is used to transform the parameters $\mathbf{A}$ and $\mathbf{B}$ to their discretized counterparts $\bar{\mathbf{A}}$ and $\bar{\mathbf{B}}$. In Mamba [Gu and Dao(2023)] and its following works [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], this is achieved with the following zero-order hold (ZOH) rule:
$\bar{\mathbf{A}}=\exp{(\mathbf{\Delta}\mathbf{A})}\mathrm{,}~{}~{}~{}\bar{\mathbf{B}}=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp{(\mathbf{\Delta}\mathbf{A})}-\mathbf{I})\cdot\mathbf{\Delta}\mathbf{B}$ (2)
Afterwards, an input sequence $\{x_{i}\}$ (for $i=1,2,...$) can be mapped to its output sequence $\{y_{i}\}$ in a similar way:
$h^{\prime}_{i}=\bar{\mathbf{A}}h_{i-1}+\bar{\mathbf{B}}x_{i}\mathrm{,}~{}~{}~{}y_{i}=\mathbf{C}h^{\prime}_{i}+\mathbf{D}x_{i}$ (3)
Mamba. Since SSMs are often used to model LTI systems, their model parameters are shared by all time steps $i$. However, as found in Mamba [Gu and Dao(2023)], such time-invariant characteristics severely limit the model’s representativity. To alleviate this problem, Mamba lifts the time-invariant constraint and makes the parameters $\mathbf{B}$, $\mathbf{C}$ and $\mathbf{\Delta}$ dependent on the input sequence $\{x_{i}\}$, a process they refer to as the selective scan, resulting in the token-dependent $\{\mathbf{B}_{i}\}$, $\{\mathbf{C}_{i}\}$ and $\{\mathbf{\Delta}_{i}\}$. Moreover, the SSM is combined with a gated MLP [Hua et al.(2022)Hua, Dai, Liu, and Le] to gain better representation ability. Specifically, the output sequence $\{y_{i}\}$ is computed from the $\{x_{i}\}$ as the following:
$x^{\prime}_{i}=\sigma\bigl{(}$ $\mathrm{DWConv}\bigl{(}\mathrm{Linear}(x_{i})\bigr{)}\bigr{)}\mathrm{,}~{}~{}~{}z_{i}=\sigma\bigl{(}\mathrm{Linear}(x_{i})\bigr{)}$ (4)
$\mathbf{B}_{i},\mathbf{C}_{i},$ $\mathbf{\Delta}_{i}=\mathrm{Linear}(x^{\prime}_{i})\mathrm{,}~{}~{}~{}\bar{\mathbf{A}}_{i},\bar{\mathbf{B}}_{i}=\mathrm{ZOH}(\mathbf{A},\mathbf{B_{i}},\mathbf{\Delta_{i}})$ (5)
$h^{\prime}_{i}=\bar{\mathbf{A}}_{i}$ $h_{i-1}+\bar{\mathbf{B}}_{i}x^{\prime}_{i}\mathrm{,}~{}~{}~{}y^{\prime}_{i}=\mathbf{C}_{i}h^{\prime}_{i}+\mathbf{D}x^{\prime}_{i}\mathrm{,}~{}~{}~{}y_{i}=y^{\prime}_{i}\odot z_{i}$ (6)
where $\sigma$ denotes the SiLU activation, and $\odot$ denotes element-wise multiply.
3.2 Overall architecture of PlainMamba
In Figure 2, we present the model architecture of PlainMamba. Our model is divided into three main components: (1) a convolutional tokenizer that transforms an input 2D image into visual tokens, (2) the main network with a series of $L$ identical PlainMamba blocks to learn visual representations, and (3) a task-specific head for downstream applications.
In more detail, the tokenizer will downsample the input image $I\in\mathbb{R}^{H_{I}\times W_{I}\times 3}$ into a list of visual tokens $x\in\mathbb{R}^{H\times W\times C}$, where $C$ is the channel number. We set the default down-sampling factor to 16, following ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]. After combining the initial visual tokens with positional embeddings [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] for retaining spatial information, the tokens undergo a series of transformations through the $L$ PlainMamba blocks, which are designed to simplify usage by maintaining the input-output shape consistency. The final stage of the architecture involves a task-specific head, which is dependent on the particular downstream application. For instance, in image classification tasks, the image tokens are globally pooled into a vector, which is then fed into a linear classification head to produce the final output.
PlainMamba distinguishes itself from existing vision transformers [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby, Liu et al.(2024a)Liu, Yang, Zhou, Xi, Yu, Yu, Liang, Shi, Zhang, Zheng, et al.] and concurrent vision Mamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] architectures in several key aspects. Firstly, it does not use any special tokens, such as the commonly used CLS token. Secondly, in contrast to approaches that adopt a hierarchical structure to manage feature resolution [Wang et al.(2022)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao, Li et al.(2022b)Li, Wu, Fan, Mangalam, Xiong, Malik, and Feichtenhofer, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], Instead, PlainMamba maintains a constant feature resolution across all blocks. This design choice considers the recent progress made in various visual foundation models [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick, Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski] where the plain non-hierarchical ViT is used rather than its hierarchical counterparts.
3.3 PlainMamba Block
The overall architecture comprises several identical PlainMamba blocks, forming the backbone for learning high-quality visual features. We present the structure of the PlainMamba block in Figure 2, in which we make several key adjustments to the original Mamba block to fully exploit the two-dimensional nature of image inputs. This adaptation is crucial for effectively transitioning from the inherently 1D processing paradigm of language models to the 2D domain of images. To this end, we introduce two novel techniques: (1) Continuous 2D Scanning and (2) Direction-Aware Updating. The first technique ensures that each visual token is always adjacent to the previous scanned token. By doing so, it mitigates positional bias and encourages a more uniform understanding of the image space, enhancing the model’s ability to learn from visual inputs. The second technique explicitly embeds the 2D relative positional information into the selective scanning process, which allows the model to better interpret the positional context of flattened visual tokens.
Figure 3: Comparison between our Continuous 2D Scanning and the selective scan orders in ViM [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu]. Our method makes sure that every scanned visual token is spatially adjacent to its predecessor, avoiding potential spatial and semantic discontinuity.
Continuous 2D Scanning. The selective scan mechanism is inherently designed for sequential data, such as text. Adapting this mechanism for 2D image data requires flattening the 2D viusal tokens into a 1D sequence to apply the State Space Model (SSM) updating rule. Prior research, e.g., VisionMamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], has demonstrated the efficacy of using multiple scanning orders to enhance model performance — such as both row-wise and column-wise scans in multiple directions. However, as shown in Figure 3 (a) and (b), in these approaches, each scanning order can only cover one type of 2D direction, e.g., left to right, causing spatial discontinuity when moving to a new row (or column). Moreover, as the parameter $\mathbf{A}$ in Equation 3 serves as a decaying term, such spatial discontinuity can also cause adjacent tokens to be decayed to different degrees, compounding the semantic discontinuity and resulting in potential performance drop.
Table 1: PlainMamba  variants. FLOPs are measured using input size 224$\times$224.
Model Depth Channels Params FLOPs PlainMamba-L1 24 192 7.3M 3.0G PlainMamba-L2 24 384 25.7M 8.1G PlainMamba-L3 36 448 50.5M 14.4G
Our Continuous 2D Scanning addresses this challenge by ensuring a scanned visual token is always adjacent (in the 2D space) to the previously scanned token. As shown in Figure 3 (c), in our approach, the visual tokens are also scanned in four distinct orders. When reaching the end of a row (or column), the next scanned token will be its adjacent, not the opposite, token in the next column (or row). Then, the scanning continues with a reversed direction until it reaches the final visual token of the image. As a consequence, our method preserves spatial and semantic continuity and avoids potential information loss when scanning non-adjacent tokens. Furthermore, in practice the model usually takes input images of the same size, meaning our method can be easily implemented and efficiently run by pre-computing the permutation indexes.
Direction-Aware Updating. As shown in Equation 3, the contribution of a token $x_{i}$ to the hidden state $h_{i}$ in the selective scan is determined by the parameter $\bar{\mathbf{B}}_{i}$, derived from $x_{i}$ itself. In language models, the sequential order naturally dictates the positional relationship between tokens, allowing the model to "remember" their relative positions. However, in our Continuous 2D Scanning, the current token can be in one of four possible directions relative to its predecessor. This challenges the model’s ability to discern the precise spatial relationship between consecutive tokens based on $\mathbf{B}_{i}$ alone. Our Direction-Aware Updating is therefore proposed to address this challenge. Drawing inspiration from the relative positional encoding mechanisms in vision transformers [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby], we employ a set of learnable parameters $\{\mathbf{\Theta}_{k}\in\mathbb{R}^{m\times 1}\}$ (for $k=1,2,...,5$), representing the four cardinal directions plus a special BEGIN direction for the initial token. These parameters are summed with the data-dependent $\mathbf{B}_{i}$ to enrich the selective scan process with directional information. Specifically, with $x_{i}$ and $z_{i}$ following Equation 3, our Direction-Aware Updating is formulated as follows:
$h^{\prime}_{k,i}=\bar{\mathbf{A}}_{i}h_{k,i-1}+(\bar{\mathbf{B}}_{i}+\bar{\mathbf{\Theta}}_{k,i})x_{i}$ (7)
$y^{\prime}_{i}=$ $\sum_{k=1}^{4}\bigl{(}\mathbf{C}_{i}h^{\prime}_{k,i}+\mathbf{D}x_{i}\bigr{)}\mathrm{,}~{}~{}~{}y_{i}=y^{\prime}_{i}\odot z_{i}$ (8)
where $k$ spans the four distinct scanning directions introduced by our Continuous 2D Scanning. Alternatively, for the initial token of each scan, we instead add the final $\bar{\mathbf{\Theta}}_{k=5}$ vector. The term $\bar{\mathbf{\Theta}}_{k,i}$ represents the discretized $\mathbf{\Theta}_{k,i}$ using $\mathbf{\Delta}_{i}$.
3.4 Model Variants of PlainMamba
As shown in Table 1, we present three different model variants of PlainMamba. Specifically, from PlainMamba-L1 to PlainMamba-L2, we scale the model width, i.e., feature channel numbers, and keep the model depth to 24. From PlainMamba-L2 to PlainMamba-L3, we scale both model width and depth. The FLOPs are measured using 224$\times$224 inputs, and we follow the official Mamba codebase to compute the FLOPs of the selective scan process.
4 Experiments
Table 2: Comparison between PlainMamba and other models on ImageNet-1K. ( denotes best epoch result.)
Model Hierarchical Params FLOPs Top-1 CNN ResNeXt101-32$\times$[Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 44M 8.0G 78.6 ResNeXt101-32$\times$[Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 88M 16.5G 79.3 RegNetY-4G [Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár] 21M 4.0G 80.0 RegNetY-8G [Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár] 39M 8.0G 81.7 ConvNeXt-T [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] 29M 4.5G 82.1 ConvNeXt-S [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] 50M 8.7G 83.1 Transformer DeiT-Tiny [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 5M 1.3G 72.2 DeiT-Small [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 22M 4.6G 79.9 DeiT-Base [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 86M 16.8G 81.8 Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 29M 4.5G 81.3 Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 50M 8.7G 83.0 PVT-Tiny [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 13M 2G 75.1 PVT-Small [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 25M 3.8G 79.8 PVT-Medium [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 44M 6.7G 81.2 Focal-Tiny [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 29M 4.9G 82.2 Focal-Small [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 51M 9.1G 83.5 State Space Modeling ViM-T [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 7M - 76.1 ViM-S [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 26M - 80.5 LocalViM-T [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 8M 1.5G 76.2 LocalViM-S [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 28M 4.8G 81.2 Mamba-ND-T [Li et al.(2024)Li, Singh, and Grover] 24M - 79.2 Mamba-ND-S [Li et al.(2024)Li, Singh, and Grover] 63M - 79.4 S4ND-ViT-B [Nguyen et al.(2022)Nguyen, Goel, Gu, Downs, Shah, Dao, Baccus, and Ré] 89M - 80.4 S4ND-ConvNeXt-T [Nguyen et al.(2022)Nguyen, Goel, Gu, Downs, Shah, Dao, Baccus, and Ré] 30M - 82.2 VMamba-T [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 22M 5.6G 82.2 VMamba-S [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 44M 11.2G 83.5 PlainMamba-L1 7M 3.0G 77.9 PlainMamba-L2 25M 8.1G 81.6 PlainMamba-L3 50M 14.4G 82.3
In the main paper, we quantitatively compare PlainMamba with previously proposed models on four visual recognition tasks: image classification, object detection, instance segmentation, and semantic segmentation. Please refer to our supplementary materials for further ablation studies.
4.1 Experiment Settings
ImageNet Classification. We build our codebase following [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], which is a commonly used training recipe. Specifically, for the ImageNet-1k experiments, we train all PlainMamba models for 300 epochs using AdamW optimizer. Following [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], we set the batch size to 2048, weight decay to 0.05, and the peak learning rate to 0.002. Cosine learning rate scheduling is used. For data augmentation, we used the commonly used recipe [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo, Ding et al.(2022)Ding, Xiao, Codella, Luo, Wang, and Yuan], which includes Mixup [Zhang et al.(2017)Zhang, Cisse, Dauphin, and Lopez-Paz], Cutmix [Yun et al.(2019)Yun, Han, Oh, Chun, Choe, and Yoo], Random erasing [Zhong et al.(2020)Zhong, Zheng, Kang, Li, and Yang] and Rand augment [Cubuk et al.(2020)Cubuk, Zoph, Shlens, and Le].
ADE20K Semantic Segmentation. We follow common practice [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo, Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang] to use UperNet [Xiao et al.(2018)Xiao, Liu, Zhou, Jiang, and Sun] as the segmentation network. Unlike XCiT [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.], we do not explicitly resize the constant resolution feature maps into multi-scale. Following [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], we train all models for 160 iterations with batch size 16 and set the default training image size to 512$\times$512.
COCO Object Detection and Instance Segmentation. Following [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], we test PlainMamba’s ability on COCO object detection and instance segmentation using both the two-stage Mask R-CNN [He et al.(2017)He, Gkioxari, Dollar, and Girshick] and the single-stage RetinaNet [Lin et al.(2017b)Lin, Goyal, Girshick, He, and Dollar]. For both models, we report the results of both 1$\times$ schedule. Following [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], we use ViTAdapter [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] to compute multi-scale features to fit the FPN network structure. We use the commonly used training settings proposed in [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] to keep a fair comparison.
4.2 Main Results
ImageNet-1K Classification. In Table 2, we report the ImageNet-1K experiment results. We compare PlainMamba with three different kinds of visual feature extractors: CNNs, vision transformers, and SSMs. In addition, the comparison includes both hierarchical and non-hierarchical models. Firstly, when comparing with SSMs, our model is doing better than the recently proposed Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and Mamba-ND [Li et al.(2024)Li, Singh, and Grover]. For example, PlainMamba-L2 achieves a 2.4% higher accuracy than Mamba-ND-T while they share a similar model size. These results validate PlainMamba’s effectiveness as a non-hierarchical SSM. Secondly, when compared with CNNs and transformers, our model achieves better performance than the non-hierarchical counterparts. For example, PlainMamba-L2 achieves 1.7% better accuracy with DeiT-Small. Moreover, PlainMamba also achieves similar performance when compared with hierarchical models. For example, when the model size is around 25M, our model achieves 0.3% better accuracy than Swin-Tiny, validating PlainMamba’s ability as a general feature extractor. On the other hand, the hierarchical VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], together with other hierarchical transformers, do achieve a better accuracy than ours. As we explained in Section 1, hierarchical models tend to perform better than non-hierarchical ones in visual recognition. As the main motivation of our work is to develop a simple Mamba architecture, a bit inferior ImageNet accuracy is acceptable.
Table 3: ADE20K semantic segmentation using UperNet. The FLOPs are computed using input size 512$\times$2048.
Backbone Hierarchical Params FLOPs mIoU CNN ResNet-50 [He et al.(2016)He, Zhang, Ren, and Sun] 67M 953G 42.1 ResNet-101 [He et al.(2016)He, Zhang, Ren, and Sun] 85M 1030G 44.0 ConvNeXt-T [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] 60M 939G 46.7 Transformer DeiT-S+MLN [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 58M 1217G 43.8 DeiT-B+MLN [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 144M 2007G 45.5 XCiT-T12/8 [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] 34M - 43.5 XCiT-S12/8 [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] 52M 1237G 46.6 XCiT-S24/8 [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] 74M 1587G 48.1 Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 60M 945G 44.5 Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 81M 1038G 47.6 Focal-Tiny [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 62M 998G 45.8 Focal-Small [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 85M 1130G 48.0 Twins-SVT-Small [Chu et al.(2021)Chu, Tian, Wang, Zhang, Ren, Wei, Xia, and Shen] 54M 912G 46.2 Twins-SVT-Small [Chu et al.(2021)Chu, Tian, Wang, Zhang, Ren, Wei, Xia, and Shen] 88M 1044G 47.7 State Space Modeling ViM-T [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 13M - 41.0 ViM-S [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 46M - 44.9 LocalVim-T [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 36M 181G 43.4 LocalVim-S [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 58M 297G 46.4 VMamba-T [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 55M 964G 47.3 VMamba-S [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 76M 1081G 49.5 PlainMamba-L1 35M 174G 44.1 PlainMamba-L2 55M 285G 46.8 PlainMamba-L3 81M 419G 49.1
ADE20K Semantic Segmentation We report our model’s ADE20K semantic segmentation performance in Table 3. Similar to the ImageNet-1k and COCO experiments, here the competing models include both hierarchical and non-hierarchical backbones in three types of visual feature extractors. The results again suggest that PlainMamba achieves the best performance among the non-hierarchical models. For example, with similar parameter amounts, PlainMamba-L2 outperforms the high-resolution (patch size of 8) XCiT-S12/8 model [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] with a much lower computation cost. Moreover, PlainMamba-L2 also outperforms the hierarchical Swin-Transformer-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], achieving better mIoU while having a lower model size and FLOPs. At the same time, PlainMamba is also doing better than the concurrent Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. For instance, PlainMamba-L2 achieves a 1.9 higher mIoU than ViM-S. This result verifies our model’s effectiveness in extracting fine-grained visual features, which is essential for the pixel-wise semantic segmentation task.
Table 4: Mask R-CNN object detection and instance segmentation on MS COCO mini-val using 1$\times$ schedule. We use ViTAdapter [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] to compute multi-scale features. FLOPs are computed using input size 1280$\times$800.
Backbone Hierarchical Params FLOPs $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ $AP^{mk}$ $AP^{mk}_{50}$ $AP^{mk}_{75}$ CNN ResNeXt101-32x4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 63M 340G 41.9 - - 37.5 - - ResNeXt101-64x4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 102M 493G 42.8 - - 38.4 - - Transformer ViT-Adapter-T [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] 29M 349G 41.1 62.5 44.3 37.5 59.7 39.9 ViT-Adapter-S [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] 49M 463G 44.7 65.8 48.3 39.9 62.5 42.8 ViT-Adapter-B [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] 131M 838G 47.0 68.2 51.4 41.8 65.1 44.9 PVT-Small [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 44M - 40.4 62.9 43.8 37.8 60.1 40.3 PVT-Medium [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 64M - 42.0 64.4 45.6 39.0 61.6 42.1 PVT-Large [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 81M - 42.9 65.0 46.6 39.5 61.9 42.5 Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 48M 264G 42.2 - - 39.1 - - Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 69M 354G 44.8 - - 40.9 - - ViL-Tiny [Zhang et al.(2021)Zhang, Dai, Yang, Xiao, Yuan, Zhang, and Gao] 26M 145G 41.4 63.5 45.0 38.1 60.3 40.8 ViL-Small [Zhang et al.(2021)Zhang, Dai, Yang, Xiao, Yuan, Zhang, and Gao] 45M 218G 44.9 67.1 49.3 41.0 64.2 44.1 ViL-Medium [Zhang et al.(2021)Zhang, Dai, Yang, Xiao, Yuan, Zhang, and Gao] 60M 293G 47.6 69.8 52.1 43.0 66.9 46.6 State Space Modeling EfficientVMamba-T [Pei et al.(2024)Pei, Huang, and Xu] 11M 60G 35.6 57.7 38.0 33.2 54.4 35.1 EfficientVMamba-S [Pei et al.(2024)Pei, Huang, and Xu] 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2 EfficientVMamba-B [Pei et al.(2024)Pei, Huang, and Xu] 53M 252G 43.7 66.2 47.9 40.2 63.3 42.9 VMamba-T [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 42M 262G 46.5 68.5 50.7 42.1 65.5 45.3 VMamba-S [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 64M 357G 48.2 69.7 52.5 43.0 66.6 46.4 PlainMamba-Adapter-L1 31M 388G 44.1 64.8 47.9 39.1 61.6 41.9 PlainMamba-Adapter-L2 53M 542G 46.0 66.9 50.1 40.6 63.8 43.6 PlainMamba-Adapter-L3 79M 696G 46.8 68.0 51.1 41.2 64.7 43.9
COCO Object Detection and Instance Segmentation. We report the results of Mask R-CNN object detection and instance segmentation in Table 4. With similar FLOPs and many fewer parameters, PlainMamba-L1 achieves 44.1 APbb and 39.1 APmk when using 1$\times$ training schedule, while Swin-Small achieves 44.8 APbb and 40.9 APmk. We also observe that hierarchical models tend to work better than non-hierarchical models. Although our model achieves lower performance than some hierarchical models, e.g., the concurrent VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], PlainMamba achieves the best performance among its non-hierarchical counterparts. For instance, when using 1$\times$ training schedule, PlainMamba achieves 3.1 higher APbb and 1.6 higher APmk than DeiT-T when they are both equipped with the ViTAdapter [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao]. These results demonstrate that PlainMamba is able to extract good local features, which is important to the object-level tasks like instance segmentation. On the other hand, we also admit that PlainMamba is performing worse than the hierarchical VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu]. We attribute such inferiority to the multi-resolution architecture of FPN-based [Lin et al.(2017a)Lin, Dollár, Girshick, He, Hariharan, and Belongie] Mask R-CNN, which is more naturally suitable to the hierarchical designs.
5 Conclusion
We present PlainMamba, a plain SSM-based model for visual recognition. Our model is conceptually simple because it uses no special tokens or hierarchical structure, making it a perfect counterpart to the widely used plain vision transformer. The results show that PlainMamba achieves superior performance to previous non-hierarchical models, including the concurrent SSM-based models, and can perform on par with the high-performing hierarchical models. We hope our model can serve as a baseline for future research in this area.
Acknowledgements
Funding for this research is provided in part by a studentship from the School of Engineering at the University of Edinburgh, the SENSE - Centre for Satellite Data in Environmental Science CDT, and an EPSRC New Investigator Award (EP/X020703/1).
Appendix A More Experiments
A.1 COCO Object Detection using RetinaNet
We report COCO RetinaNet object detection in Table 5. Similar to Mask R-CNN, whose results are reported in the main paper, PlainMamba also performs well with the single-stage RetinaNet object detector. For example, with only half the model size and similar FLOPs, PlainMamba-L1 achieves 0.2 higher AP than Swin-Tiny.
Table 5: RetinaNet object detection on MS COCO mini-val with 1$\times$ schedule. FLOPs are computed using input size 1280$\times$800.
Backbone Hierarchical Params FLOPs $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ CNN ResNeXt101-32x4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 56M 319G 39.9 - - ResNeXt101-64×4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 95M 413G 41.0 - - Transformer Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 38M 245G 41.5 - - Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 60M 335G 44.5 - - Focal-Tiny [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 39M 265G 43.7 - - Focal-Small [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 62M 367G 45.6 - - PVT-Small [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 34M - 40.4 61.3 43.0 PVT-Medium [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 54M - 41.9 63.1 44.3 PVT-Large [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 71M - 42.6 63.7 45.4 State Space Modeling EfficientVMamba-T [Pei et al.(2024)Pei, Huang, and Xu] 13M - 37.5 57.8 39.6 EfficientVMamba-S [Pei et al.(2024)Pei, Huang, and Xu] 19M - 39.1 60.3 41.2 EfficientVMamba-B [Pei et al.(2024)Pei, Huang, and Xu] 44M - 42.8 63.9 45.8 PlainMamba-Adapter-L1 19M 250G 41.7 62.1 44.4 PlainMamba-Adapter-L2 40M 392G 43.9 64.9 47.0 PlainMamba-Adapter-L3 67M 478G 44.8 66.0 47.9
A.2 Ablation Studies and Discussions
Setting: Here, we conduct ablation studies to test our model designs and to gain a deeper understanding of the proposed method. We use our L1 model, with less than 10M parameters, for most experiments. The models are all pre-trained on ImageNet-1K following the same training settings described in the main paper.
Table 6: Ablation study of model depth v.s. width on ImageNet-1K..
     Depth Width Params FLOPs Top-1      6 376 7.3 M 2.5 G 74.6      12 272 7.5 M 2.7 G 76.8      24 192 7.3 M 3.0 G 77.9      36 156 7.2 M 3.3 G 77.9
Depth v.s. Width When designing neural architectures for a given parameter count, it’s usually important to find a good balance between the network’s depth, i.e., the number of layers, and its width, i.e., the feature dimensions. While this problem was studied for existing architectures [Wu et al.(2019)Wu, Shen, and Van Den Hengel, Xue et al.(2022)Xue, Shi, Wei, Lou, Liu, and You], it is still unclear whether the previous conclusions are applicable to vision SSMs. In Table 6, we study the depth and width trade-off of the proposed PlainMamba. Firstly, the results show that deeper models tend to perform better than shallow ones. For example, when the parameter count is around 7.4M, the 12-layer model achieves 2.2% higher ImageNet top-1 accuracy than the 6-layer counterparts, and the 24-layer model is further 1.1% higher than the 12-layer one. However, when we further increase the depth to 36 while reducing the width accordingly, the top-1 accuracy remains similar. On the other hand, we also notice that deeper models are less efficient than shallower but wider models. For instance, the 24-layer model is 0.3G FLOPs higher than the 12-layer model. These results suggest the necessity of a good balance between network depth and width.
Figure 4: Efficiency comparison between PlainMamba and DeiT. We modify the DeiT-Tiny model by changing its channel number to 224, resulting in a similar-size model (7.4M) to PlainMamba-L1. The peak memory is measured using a batch size of 1.
Table 7: Ablation study of model depth v.s. width on ImageNet-1K.
Method Params FLOPs Top-1 VisionMamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 7.8M 1.3G 74.4 VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 7.3M 3.0G 77.1 Ours 7.3M 3.0G 77.9
PlainMamba Block Design. Here, we test different designs of the PlainMamba block by comparing it with the block designs in Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu]. For a fair comparison, we use the same model depth and width settings for all designs and train all models with the same training recipe. We also remove the CLS tokens from the Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] block and use the global averaging pooling as an alternative. We report the results in Table 7. We can see that our design achieves the best results. Specifically, Vision Mamba only achieves a 74.4% ImageNet accuracy, which is 3.5% lower than ours. We also notice that the model with a Vision Mamba block is inferior to the original Vision Mamba model, which is caused by the removal of the CLS token. These results suggest that our model still retains its ability when the CLS token is absent. Also, our design performs better than the VMamba block [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] with a 0.8% accuracy advantage, indicating that the improvements come from our proposed Continuous 2D Scanning and Direction-aware Updating, which validate the effectiveness of our proposed techniques in adapting SSM for 2D images.
Table 8: Comparison of decomposed FLOPs between DeiT and PlainMamba.
Resolution Part DeiT-C224 PlainMamba-L1 128$\times$128 Token Mixing 0.18G 0.34G Channel Mixing 0.31G 0.33G Others 0.01G 0.30G 4096$\times$4096 Token Mixing 23244G 350G Channel Mixing 315G 348G Others 12G 311G
Efficiency Comparison with ViT One particular advantage of SSMs, e.g., Mamba, is their ability to capture global information while maintaining efficiency. in Figure 4, we compare the PlainMamba’s efficiency with the vision transformer. Specifically, to ensure a fair comparison, we create a DeiT model with channel numbers of 224, resulting in a model with 7.4M parameters, which is used to compare with PlainMamba-L1. Specifically, we compare the model FLOPs and the peak inference memory using inputs of different sizes. The results show that our model is able to keep the computation cost low when the input size is scaled up to high resolutions, e.g., 4096$\times$4096. However, DeiT’s FLOPs and memory consumption increase rapidly when using such high-resolution inputs. On the other hand, we also notice that our model’s efficiency is inferior to the similar-sized DeiT when using low-resolution images, e.g., 128$\times$128. To further investigate such a difference in their efficiency, we decompose their FLOPs into three parts [Tolstikhin et al.(2021)Tolstikhin, Houlsby, Kolesnikov, Beyer, Zhai, Unterthiner, Yung, Steiner, Keysers, Uszkoreit, et al.]: 1) token mixing, 2) channel mixing, and 3) others. Specifically, token mixing refers to the multi-head attention part in DeiT and the selective scanning part in PlainMamba, and channel mixing refers to the feed-forward network in DeiT and the input & output projection in PlainMamba. We report the results in Table 8. These suggest that PlainMamba’s FLOPs are evenly distributed across the three parts in low and high resolutions. On the contrary, when using 128$\times$128 inputs, DeiT’s FLOPs are dominated by channel-mixing and the other two parts are negligible. However, because of the quadratic complexity of self-attention operation, DeiT’s FLOPs in token mixing grow to 23T when using 4096$\times$4096 inputs, 23 times more expensive than PlainMamba. These results verify PlainMamba’s high efficiency for high-resolution inputs.